# Scene-Level Appearance Transfer with Semantic Correspondences

### v1

@@ Abstract
We introduce ReStyle3D, a novel framework for scene-level appearance transfer from a single style image to a real-world scene represented by multiple views. The method combines explicit semantic correspondences with multi-view consistency to achieve precise and coherent stylization. Unlike conventional stylization methods that apply a reference style globally, ReStyle3D uses open-vocabulary segmentation to establish dense, instance-level correspondences between the style and real-world images. This ensures that each object is stylized with semantically matched textures. ReStyle3D first transfers the style to a single view using a training-free semantic-attention mechanism in a diffusion model. It then lifts the stylization to additional views via a learned warp-and-refine network guided by monocular depth and pixel-wise correspondences. Experiments show that ReStyle3D consistently outperforms prior methods in structure preservation, perceptual style similarity, and multi-view coherence. User studies further validate its ability to produce photo-realistic, semantically faithful results. Our code, pretrained models, and dataset will be publicly released, to support new applications in interior design, virtual staging, and 3D-consistent stylization.
Figure 1: Given an interior design image (style image) and a 3D scene captured by video or multi-view images, ReStyle3D first transfers the appearance based on semantic correspondences to a single view, then lifts the stylization to multiple viewpoints using 3D-aware style lifting, achieving multi-view consistent appearance transfer with fine-grained details. Project page: restyle3d.github.io.
1 Introduction
These authors contributed equally to this work.
@@ 1 Introduction
Generative diffusion models have recently spurred significant advances in image stylization and broader generative applications, enabling the seamless synthesis or editing of images with remarkable visual fidelity. While existing image stylization approaches [7, 30] often excel at transferring well-known artistic styles (e.g., Van Gogh paintings) onto photographs, they fall short when it comes to practical and realistic style applications, such as virtual staging or professional interior decoration, where transferring the style of one image (style image) to another (source image) entails transferring the individual appearance of objects (Fig. 1).
@@
These methods tend to treat the style image globally, ignoring the semantic correspondence between individual objects or regions in the images. This coarsely aligned stylization not only misrepresents object appearances but also fails to adapt fine-grained textures to semantically matched regions (e.g., transferring couch textures only to couches). This is crucial for real-world use cases where style is defined by the unique characteristics (e.g., color, material, shape) of design elements (i.e., furniture, decor, lighting, and accessories) that give it its signature look [44]. Another line of work pursues semantic correspondence for transferring object appearances [5, 70]. While these methods show promise in aligning single objects or small regions via deep feature matching, they typically operate at low spatial resolutions (often $64\times 64$) and therefore struggle to handle complex scenes with strong perspective and multiple object instances. Extending them to scene-level stylization remains a challenging problem due to both semantic and geometric complexity.
@@
Moreover, when a scene is represented by multiple images (e.g., for larger coverage), ensuring multi-view consistency in scene-level appearance transfer further complicates the task. Existing multi-view image editing methods [45, 35, 16, 10] commonly require known camera poses and an existing 3D scene representation (e.g., a neural radiance field [39] or 3D Gaussian splatting [28]), which needs a dense set of input views and considerable compute time. These methods struggle with sparse or casually captured views, and their specialized 3D pipelines hinder plug-and-play use. A pixel-space approach preserving geometric cues without heavy 3D modeling is preferable but remains under explored. We propose ReStyle3D, a novel framework for scene-level appearance transfer that cobines semantic correspondence and multi-view consistency, addressing limitations of 2D stylization and 3D-based editing methods. Our key insight is that the inherent but implicit semantic correspondences from pretrained diffusion models or vision transformers (e.g., StableDiffusion [49] and DINO [4, 42]) are insufficient for fine-grained, scene-level appearance transfer, especially when different objects or viewpoints are involved. We tackle this by explicitly matching open-vocabulary panoptic segmentation predictions between the style and source images, while ensuring that unmatched parts of the scene still receive a global style harmonization. This open-vocabulary labeling (with no predefined semantic categories) helps us robustly align semantically corresponding regions even in cluttered indoor scenes. By integrating these explicit correspondences into the attention mechanism of a diffusion process, we achieve more accurate and flexible stylization of multi-object scenes.
@@
To further ensure 3D awareness and view-to-view consistency, we adopt a two-stage pipeline. First, we achieve training-free semantic appearance transfer in a single view by injecting our correspondence-informed attention into a pretrained diffusion model. Second, a warp-and-refine diffusion network that efficiently propagates the stylized appearance to additional views in an auto-regressive manner, guided by monocular depth and pixel-level optical flows. Our method does not require explicit pose or 3D modeling, and we show that the final stylized frames are fully compatible with off-the-shelf 3D reconstruction tools, enabling complete 3D visualizations and consistent multi-view stylization with minimal overhead. In summary, our contributions are as follows:
@@
  • We introduce SceneTransfer, a new task of transferring multi-object appearance from a single style image to a 3D scene captured in multi-view images.
@@
  • We propose ReStyle3D, a two-stage pipeline that (i) adapts a pretrained diffusion model with semantic attention for instance-level stylization, and (ii) trains a warp-and-refine novel-view synthesis module to propagate the style across all views, maintaining global consistency.
@@ 2 Related Work
aims to transfer artistic styles to images while preserving structural content. Early CNN-based methods [17, 24, 11] laid the groundwork by capturing style and content representations. With the advent of diffusion models [23, 49], recent approaches leverage pretrained architectures and textual guidance for high-quality stylization [7, 59, 31, 14, 65, 30, 73]. InST [73] employs textual inversion to encode styles in dedicated text embeddings, achieving flexible transfer. StyleDiffusion [31] further refines style-content separation through a CLIP-based disentanglement loss applied during fine-tuning. StyleID [7] adapts self-attention in pretrained diffusion models to incorporate artistic styles without additional training. While these methods produce compelling results, they focus on overall style transfer without explicitly modeling semantic correspondences. In contrast, we attempt to inject explicit semantic matching in stylization, thereby enabling precise style transfer according to semantically matching regions.
@@
Foundational works and recent innovations have shaped the evolution of semantic correspondence. SIFT-Flow [33] pioneered dense image alignment with handcrafted SIFT descriptors [37]. Self-supervised vision transformers like DINO [4] and DINO-V2 [42, 8] improved feature representation for semantic matching without labeled data [57, 56]. Recent methods, such as  [70, 19], DIFT [53], cross-image-attention [1] and [18], integrate diffusion models with these transformers, achieving superior zero-shot correspondence. Techniques like Deep Functional Maps [5] further refine correspondences by enforcing global structural consistency, demonstrating the potential of advanced representations in addressing correspondence challenges. The development of these techniques enables the extraction of semantic correspondences using intermediate representations.
@@
The attention modules in pretrained diffusion models are essential in controlling the generated content, allowing various image editing tasks through attention mask manipulation. Prompt-to-Prompt [20] pioneered text-based local editing by manipulating cross-attention between text prompts and image regions. Similarly, Plug-and-play [58] leverages the original image’s spatial features and self-attention maps to preserve spatial layout while generating text-guided edited images. Epstein et al. [12] introduced Diffusion Self-Guidance, a zero-shot approach that leverages internal representations for fine-grained control over object attributes. While these methods focus on text-to-image attention control, recent works like Generative Rendering [3] explore cross-image attention by injecting 4D correspondences from meshes into attention for stylized video generation. In contrast, we propose a direct image-to-image semantic attention mechanism that transfers appearances across all semantic categories simultaneously through explicit correspondence masks, enabling efficient and accurate scene-level stylization without text prompts or 3D priors.
Diffusion-based Novel-View Synthesis
of 3D scenes typically requires inferring and synthesizing new regions that are either unobserved or occluded in the original viewpoint. A paradigm in prior work [62, 47, 32, 29] is the warp-and-refine approach: estimate a depth map from the input image, warp the image to the desired viewpoint, and then fill in occluded or missing areas through a learned refinement stage. More recent research [48, 68, 26] avoids explicit depth-based warping by directly training generative models that handle view synthesis in a single feed-forward pass. Another line of work [43, 54, 6, 51, 40, 2, 55, 69, 52, 9, 50] integrates diffusion models such as StableDiffusion [49], making it possible to extrapolate plausible new views that are far from the input image for in-the-wild contents. ReconX [34] and ViewCrafter [69] both harness powerful video diffusion models combined with coarse 3D structure guidance to mitigate sparse-view ambiguities, achieving improved 3D consistency for novel-view synthesis. Motivated by recent success in the warp-and-refine paradigm [50], we adopt a similar strategy but with a focus on style lifting, incorporating historical frames through adaptive blending to consistently propagate our style transfers across multiple views.
3 ReStyle3D
@@ 3.1 Preliminaries
progressively add noise to an image $\mathbf{I}_{0}$ sampled from a data distribution $p_{\mathrm{data}}(\mathbf{I})$, transforming it into Gaussian noise $\mathbf{I}_{T}$ over $T$ steps, following a variance schedule $\{\alpha_{t}\}_{t=1}^{T}$:
$p(~{}\mathbf{I}_{t}~{}|~{}\mathbf{I}_{0}~{})=\mathcal{N}(~{}\mathbf{I}_{t};~{}\sqrt{\alpha_{t}}~{}\mathbf{I}_{0},1-\alpha_{t}\mathbf{I}~{}),$ (1)
where $\mathbf{I}_{t}$ represents the noisy image at timestep $t$. The reverse process is performed by a denoising model $\epsilon_{\theta}(\cdot)$ that gradually removes noise from $\mathbf{I}_{t}$ to obtain cleaner $\mathbf{I}_{t-1}$. Here $\theta$ is the learnable parameters of the denoising model. During training, the denoising model is trained to remove noise following the objective function [23]:
$\mathcal{L}=\mathbb{E}_{\mathbf{I}_{0},t\sim\mathcal{U}(T),\epsilon\sim\mathcal{N}(0,I)}||\hat{\epsilon}_{\theta}-\epsilon||_{2}^{2},$ (2)
where $\hat{\epsilon}_{\theta}=\hat{\epsilon}_{\theta}(\mathbf{I}_{t},t,c)$, and $c$ is an optional input condition such as text, image mask, or depth information. At inference stage, a clean image $\mathbf{I}:=\mathbf{I}_{0}$ is reconstructed from a randomly sampled Gaussian noise $\mathbf{I}_{T}\sim\mathcal{N}(0,I)$ through an iterative noise-removal process. The cornerstone of modern image-based diffusion models is the latent diffusion model [49] (LDM), where the diffusion process is brought to the latent space [13] of a variational autoencoder (VAE). This approach is significantly more efficient compared to working directly in the pixel space.
Figure 2: Semantic Appearance Transfer. The style and source images are first noised back to step $T$ using DDPM inversion [25]. During the generation of the stylized output, the extended self-attention layer transfers style information from the style to the output latent. This process is further guided by a semantic matching mask, which allows for precise control.
@@
are fundamental building blocks in LDM. Given an intermediate feature map $F\in\mathbb{R}^{L\times d_{h}}$, where $L$ denotes the feature length and $d_{h}$ represents the feature dimension, the attention layer captures the interactions between all pairs of features through query-key-value operations:
$\begin{split}\phi&=\text{softmax}\left(\frac{Q^{\prime}\cdot K^{\prime T}}{\sqrt{d_{h}}}\right)\cdot V^{\prime}\\ Q^{\prime}&=Q\cdot W_{q},\quad K^{\prime}=K\cdot W_{k},\quad V^{\prime}=V\cdot W_{v},\end{split}$ (3)
where $\phi$ is the updated feature map, $Q^{\prime},K^{\prime}$, and $V^{\prime}$ are linearly projected representations of the inputs via $W_{q}$, $W_{k}$, and $W_{v}$, respectively. In self-attention, the key, query, and value originate from the same feature map, enabling context exchange within the same domain. For cross attention, the key and value come from a different source, facilitating information exchange across domains. In ReStyle3D, we tailor the self-attention layers specifically for semantic appearance transfer and keep the cross-attention unchanged.
3.2 Appearance Transfer via Semantic Matching
To transfer the appearance of $\mathbf{I}_{style}$ to $\mathbf{I}_{src}$, prior attempts also employing diffusion models [1, 70, 5] have primarily focused on single objects, and struggle with scene-level transfer involving multiple instances. Our key observation is that the implicit semantic correspondences in foundation models [49, 42] are insufficient for more complex multi-instance semantic matching. To address this limitation, ReStyle3D explicitly establishes and leverages semantic correspondences throughout the transfer process.

### v2

@@ Abstract.
We introduce ReStyle3D, a novel framework for scene-level appearance transfer from a single style image to a real-world scene represented by multiple views. The method combines explicit semantic correspondences with multi-view consistency to achieve precise and coherent stylization. Unlike conventional stylization methods that apply a reference style globally, ReStyle3D uses open-vocabulary segmentation to establish dense, instance-level correspondences between the style and real-world images. This ensures that each object is stylized with semantically matched textures. ReStyle3D first transfers the style to a single view using a training-free semantic-attention mechanism in a diffusion model. It then lifts the stylization to additional views via a learned warp-and-refine network guided by monocular depth and pixel-wise correspondences. Experiments show that ReStyle3D consistently outperforms prior methods in structure preservation, perceptual style similarity, and multi-view coherence. User studies further validate its ability to produce photo-realistic, semantically faithful results. Our code, pretrained models, and dataset will be publicly released, to support new applications in interior design, virtual staging, and 3D-consistent stylization. Project page and code at https://restyle3d.github.io/.
Appearance Transfer, Image Stylization, Diffusion Model, Semantic Correspondences.
submissionid: 66journalyear: 2025copyright: acmlicensedconference: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ; August 10–14, 2025; Vancouver, BC, Canadabooktitle: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers (SIGGRAPH Conference Papers ’25), August 10–14, 2025, Vancouver, BC, Canadadoi: 10.1145/3721238.3730655isbn: 979-8-4007-1540-2/2025/08ccs: Computing methodologies Computational photographyccs: Computing methodologies Image processingccs: Computing methodologies Computer vision
1. Introduction
@@
Generative diffusion models have recently spurred significant advances in image stylization and broader generative applications, enabling the seamless synthesis or editing of images with remarkable visual fidelity. While existing image stylization approaches (Chung et al., 2024; Li, 2024) often excel at transferring well-known artistic styles (e.g., Van Gogh paintings) onto photographs, they fall short when it comes to practical and realistic style applications, such as virtual staging or professional interior decoration, where transferring the style of one image (style image) to another (source image) entails transferring the individual appearance of objects (see Fig. 1).
@@
These methods tend to treat the style image globally, ignoring the semantic correspondence between individual objects or regions in the images. This coarsely aligned stylization not only misrepresents object appearances but also fails to adapt fine-grained textures to semantically matched regions (e.g., transferring couch textures only to couches). This is crucial for real-world use cases where style is defined by the unique characteristics (e.g., color, material, shape) of design elements (i.e., furniture, decor, lighting, and accessories) that give it its signature look (Park and Hyun, 2022). Another line of work pursues semantic correspondence for transferring object appearances (Cheng et al., 2024; Zhang et al., 2023a). While these methods show promise in aligning single objects or small regions via deep feature matching, they typically operate at low spatial resolutions (often $64\times 64$) and therefore struggle to handle complex scenes with strong perspective and multiple object instances. Extending them to scene-level stylization remains a challenging problem due to both semantic and geometric complexity.
@@
Moreover, when a scene is represented by multiple images (e.g., for larger coverage), ensuring multi-view consistency in scene-level appearance transfer further complicates the task. Existing multi-view editing methods (Patashnik et al., 2024; Liu et al., 2024b; Fujiwara et al., 2024; Dong and Wang, 2023) commonly require known camera poses and an existing 3D scene representation (e.g., a neural radiance field (Mildenhall et al., 2021) or 3D Gaussian splatting (Kerbl et al., 2023)), which needs a dense set of input views and considerable compute time. These methods struggle with sparse or casually captured views, and their specialized 3D pipelines hinder plug-and-play use. A pixel-space approach preserving geometric cues without heavy 3D modeling is preferable but remains under explored. We propose ReStyle3D, a novel framework for scene-level appearance transfer that combines semantic correspondence and multi-view consistency, addressing limitations of 2D stylization and 3D-based editing methods. Our key insight is that the inherent but implicit semantic correspondences from pretrained diffusion models or vision transformers (e.g., StableDiffusion (Rombach et al., 2022) and DINO (Caron et al., 2021; Oquab et al., 2024)) are insufficient for fine-grained, scene-level appearance transfer, especially when different objects or viewpoints are involved. We tackle this by explicitly matching open-vocabulary panoptic segmentation predictions between the style and source images, while ensuring that unmatched parts of the scene still receive a global style harmonization. This open-vocabulary labeling (with no predefined semantic categories) helps us robustly align semantically corresponding regions even in cluttered indoor scenes. By integrating these explicit correspondences into the attention mechanism of a diffusion process, we achieve more accurate and flexible stylization of multi-object scenes.
@@
To further ensure 3D awareness and view-to-view consistency, we adopt a two-stage pipeline. First, we achieve training-free semantic appearance transfer in a single view by injecting our correspondence-informed attention into a pretrained diffusion model. Second, a warp-and-refine diffusion network that efficiently propagates the stylized appearance to additional views in an auto-regressive manner, guided by monocular depth and pixel-level optical flows. Our method does not require explicit pose or 3D modeling, and we show that the final stylized frames are fully compatible with off-the-shelf 3D reconstruction tools, enabling complete 3D visualizations and consistent multi-view stylization with minimal overhead.
In summary, our contributions are as follows:
@@
  • We introduce SceneTransfer, a new task of compositionally transferring multi-object appearance from a single style image to a 3D scene captured in multi-view images or video.
@@
  • We propose ReStyle3D, a two-stage pipeline that (i) repurposes a pretrained diffusion model with semantic attention for instance-level stylization, and (ii) trains a warp-and-refine novel-view synthesis module to propagate the style across all views, maintaining global consistency.
@@ 2. Related Work
The goal is to transfer artistic styles to images while preserving structural content. Early CNN-based methods (Gatys et al., 2016; Huang and Belongie, 2017; Dumoulin et al., 2017) laid the groundwork by capturing style and content representations. With the advent of diffusion models (Ho et al., 2020; Rombach et al., 2022), recent approaches leverage pretrained architectures and textual guidance for high-quality stylization (Chung et al., 2024; Šubrtová et al., 2023; Li et al., 2023; Everaert et al., 2023; Yang et al., 2023; Li, 2024; Zhang et al., 2023b). InST (Zhang et al., 2023b) employs textual inversion to encode styles in dedicated text embeddings, achieving flexible transfer. StyleDiffusion (Li et al., 2023) further refines style-content separation through a CLIP-based disentanglement loss applied during fine-tuning. StyleID (Chung et al., 2024) adapts self-attention in pretrained diffusion models to incorporate artistic styles without additional training. While these methods produce compelling results, they focus on overall style transfer without explicitly modeling semantic correspondences. In contrast, we attempt to inject semantic matching in stylization, thereby enabling precise style transfer according to semantically matching regions.
@@
Foundational works and recent innovations have shaped the evolution of semantic correspondence. SIFT-Flow (Liu et al., 2011) pioneered dense image alignment with handcrafted SIFT descriptors (Lowe, 2004). Self-supervised vision transformers like DINO (Caron et al., 2021) and DINO-V2 (Oquab et al., 2024; Darcet et al., 2024) improved feature representation for semantic matching without labeled data (Tumanyan et al., 2023a, 2022). Recent methods, such as  (Zhang et al., 2023a; Hedlin et al., 2023), DIFT (Tang et al., 2023), cross-image-attention (Alaluf et al., 2024), and (Go et al., 2024), integrate diffusion models with these transformers, achieving superior zero-shot correspondence. Techniques like Deep Functional Maps (Cheng et al., 2024) further refine correspondences by enforcing global structural consistency, demonstrating the potential of advanced representations in addressing correspondence challenges. The development of these techniques enables the extraction of semantic correspondences using intermediate representations.
@@
The attention modules in pretrained diffusion models are essential in controlling the generated content, allowing various image editing tasks through attention mask manipulation. Prompt-to-Prompt (Hertz et al., 2022) pioneered text-based local editing by manipulating cross-attention between text prompts and image regions. Similarly, Plug-and-play (Tumanyan et al., 2023b) leverages the original image’s spatial features and self-attention maps to preserve spatial layout while generating text-guided edited images. Epstein et al. (2023) introduced Diffusion Self-Guidance, a zero-shot approach that leverages internal representations for fine-grained control over object attributes. While these methods focus on text-to-image attention control, recent works like Generative Rendering (Cai et al., 2024) explore cross-image attention by injecting 4D correspondences from meshes into attention for stylized video generation. MasaCtrl (Cao et al., 2023) proposed text-based non-rigid image synthesis by injecting attention masks between text and image. In contrast, we propose a direct image-to-image semantic attention mechanism that transfers appearances across all semantic categories simultaneously through explicit correspondence masks, enabling efficient and accurate scene-level stylization without text prompts or 3D priors.
Diffusion-based Novel-View Synthesis (NVS)
NVS of general scenes typically requires inferring and synthesizing new regions that are either unobserved or occluded in the original viewpoint. A common strategy in prior work (Wiles et al., 2020; Rockwell et al., 2021; Liu et al., 2021; Koh et al., 2022) is to follow a warp-and-refine approach: estimate a depth map from the input image, warp the image to the desired viewpoint, and then fill in occluded or missing areas through a learned refinement stage. More recent research (Rombach et al., 2021; Yu et al., 2023; Jin et al., 2024) avoids explicit depth-based warping by directly training generative models that handle view synthesis in a single feed-forward pass. StoryDiffusion (Zhou et al., 2024) proposes consistent self-attention to boost long-term consistency. Another line of work (Ouyang et al., 2023; Chung et al., 2023; Shriram et al., 2025; Cai et al., 2023; Tseng et al., 2023; Yu et al., 2024; Sun et al., 2024; Deng et al., 2024; Seo et al., 2024) integrates diffusion models such as StableDiffusion (Rombach et al., 2022), making it possible to extrapolate plausible new views that are far from the input image for in-the-wild contents. ReconX (Liu et al., 2024a) and ViewCrafter (Yu et al., 2024) both harness powerful video diffusion models combined with coarse 3D structure guidance to mitigate sparse-view ambiguities, achieving improved 3D consistency for novel-view synthesis. Motivated by recent success in the warp-and-refine paradigm (Seo et al., 2024), we adopt a similar strategy but with a focus on style lifting, incorporating historical frames through adaptive blending to consistently propagate our style transfers across multiple views.
3. ReStyle3D
@@ 3.1. Preliminaries
Diffusion processes progressively add noise to an image $\mathbf{I}_{0}$ sampled from a data distribution $p_{\mathrm{data}}(\mathbf{I})$, transforming it into Gaussian noise $\mathbf{I}_{T}$ over $T$ steps, following a variance schedule $\{\alpha_{t}\}_{t=1}^{T}$:
(1) $p(~{}\mathbf{I}_{t}~{}|~{}\mathbf{I}_{0}~{})=\mathcal{N}(~{}\mathbf{I}_{t};~{}\sqrt{\alpha_{t}}~{}\mathbf{I}_{0},1-\alpha_{t}\mathbf{I}~{}),$
where $\mathbf{I}_{t}$ represents the noisy image at timestep $t$. The reverse process is performed by a denoising model $\epsilon_{\theta}(\cdot)$ that gradually removes noise from $\mathbf{I}_{t}$ to obtain cleaner $\mathbf{I}_{t-1}$. Here $\theta$ is the learnable parameters of the denoising model. During training, the denoising model is trained to remove noise following the objective function (2020):
(2) $\mathcal{L}=\mathbb{E}_{\mathbf{I}_{0},t\sim\mathcal{U}(T),\epsilon\sim\mathcal{N}(0,I)}||\hat{\epsilon}_{\theta}-\epsilon||_{2}^{2},$
where $\hat{\epsilon}_{\theta}=\hat{\epsilon}_{\theta}(\mathbf{I}_{t},t,c)$, and $c$ is an optional input condition such as text, image mask, or depth information. At inference stage, a clean image $\mathbf{I}:=\mathbf{I}_{0}$ is reconstructed from a randomly sampled Gaussian noise $\mathbf{I}_{T}\sim\mathcal{N}(0,I)$ through an iterative noise-removal process. The cornerstone of modern image-based diffusion models is the latent diffusion model (Rombach et al., 2022) (LDM), where the diffusion process is brought to the latent space (Esser et al., 2021) of a variational autoencoder (VAE). This approach is significantly more efficient compared to working directly in the pixel space.
Figure 2. Semantic Appearance Transfer. The style and source images are first noised back to step $T$ using DDPM inversion (2024). During the generation of the stylized output, the extended self-attention layer transfers style information from the style to the output latent. This process is further guided by a semantic matching mask, which allows for precise control.
@@
Attention layers are fundamental building blocks in LDM. Given an intermediate feature map $F\in\mathbb{R}^{L\times d_{h}}$, where $L$ denotes the feature length and $d_{h}$ represents the feature dimension, the attention layer captures the interactions between all pairs of features through query-key-value operations:
(3) $\begin{split}\phi&=\text{softmax}\left(\frac{Q^{\prime}\cdot K^{\prime T}}{\sqrt{d_{h}}}\right)\cdot V^{\prime}\\ Q^{\prime}&=Q\cdot W_{q},\quad K^{\prime}=K\cdot W_{k},\quad V^{\prime}=V\cdot W_{v},\end{split}$
where $\phi$ is the updated feature map, $Q^{\prime},K^{\prime}$, and $V^{\prime}$ are linearly projected representations of the inputs via $W_{q}$, $W_{k}$, and $W_{v}$, respectively. In self-attention, the key, query, and value originate from the same feature map, enabling context exchange within the same domain. For cross attention, the key and value come from a different source, facilitating information exchange across domains. In ReStyle3D, we tailor the self-attention layers specifically for semantic appearance transfer, while leaving the cross-attention layers unchanged.
3.2. Appearance Transfer via Semantic Matching
To transfer the appearance of $\mathbf{I}_{style}$ to $\mathbf{I}_{src}$, prior attempts also employing diffusion models (Alaluf et al., 2024; Zhang et al., 2023a; Cheng et al., 2024) have primarily focused on single objects, and struggle with scene-level transfer involving multiple instances. Our key observation is that the implicit semantic correspondences in foundation models (Rombach et al., 2022; Oquab et al., 2024) are insufficient for more complex multi-instance semantic matching. To address this limitation, ReStyle3D explicitly establishes and leverages semantic correspondences throughout the transfer process.